RC Update and Run Plan for Weekend March 14-15, 2020

From HallCWiki
Jump to: navigation, search

RC and Meeting Information

Current RC: Jay Benesch benesch@jlab.org (Filling in for Whit starting Sunday afternoon).


RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

General updates:

  • Hall is in restricted access.
  • Let the RC, Jay Benesch, and Walter Kellner know about any non-pivot work you'd like to take place during the March 13-20 down in Hall C. (Please CC Brad as well.)
  • Jay Benesch now has the RC phone
  • The work items are currently:
    1. A 3He compressor checkout for the x > 1 experiment next year (Meekins/RadCon);
    2. SHMS dipole PS relay replacement followed by testing (Lassiter/Segal);
    3. Replace a gauge in the SHMS detector hut and put a lead blanket around it (Kenyon)
    4. SOLID detector test the scope of which is defined in the HClist http://devweb.acc.jlab.org/CSUEApps/hclist/task/1809 (Malace et al).

Current Run Plan:

Monday

  • Survey and alignment
       Confirm target and beamline laser alignment
       Support setup for Field and Compass measurements (start in Swing?)
  • continue Solid prototype installation

Tuesday - Wednesday

  • Reinstall beamline window spools, pump down beamlines when ready (install protective window guards)
  • Possible pointing Surveys for SHMS, HMS (TBD)
  • Field and Compass measurement continue as needed
  • Install reference cell, leak check, etc.
  • Install target cell and ladder as soon as possible
  • Recommission target and start spinning up ASAP (through SWING and OWL as needed)

Notes for the Future

  • At d2n startup, it would be a good idea to quantitatively check the dispersion at the target. I (D.Mack) don't think we need a tight spec, but if it's as sloppy as 1 cm/% (hence a decent spectrometer), then we're more at risk for things like how much the beam moves when the locks are off, or scraping from an off-energy tail on the beam.

http://opsweb.acc.jlab.org/CSUEApps/atlis/atlis.php?load=Task&task_id=20434 is an un-approved plan to revert to the 6 GeV style optics with one dispersion peak and more effective horizontal dispersion control. Vertical dispersion control: no such animal on fifth pass due to spreader/recombiner design.


  • 'Filler' runs for next spin up / target-down period (Don't want to forget about this!):
    • DONE 1 hour run with target out for measure of the background just from the beryllium windows. (Needs SHMS, btw.)

Pending Issues

  • Be sure to restart main alarm handlers and turn up speaker volume before operating magnets! [1]
    Be sure to restart target alarm handler and turn up speaker volume before operating target! [2]
    pNMR issues
    software 
    Software/driver issue needs to be resolved. Checkout during next non-production down time (Moller or SHMS fix)
    EPR problems 
    [3] saw some progress on Tuesday [4], but issues remain.
    Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares).
    It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
    Kepco power supply 
    PS commissioning/installation (Todd, Bill) (Bill got a filter design from the vendor and has ordered parts.)
    New filter (for 180 Hz) to install for current mode operation fix. Can always go to voltage mode if it doesn't work.
    NMR Calibration Constants 
    Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted
    [5] [6] [7] [8] [9] (Bill, Junhao, Arun)
    Beam trips 
    Persistent BLM and ion chamber trips
    • General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
    • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
    • Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
    USB bridge 
    Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    • Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
    • Ensure issues on Sat, March 14 have been corrected [10]
    IOC issues 
    Green now: Rebooted 3/2/20 with Dave Gaskell assistance. OPs wished to reboot hung iochc10 [11], noted in December as well [12]
    This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.

Opportunistic Access Jobs

None at the moment.



Bluejeans remote connection information

 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920


Useful Links